$SPAD/src/input summation.input

Timothy Daly

December 3, 2016

Abstract

The output of the sum and summation functions is ambiguous. The mathml.spad code has been changed to fix this but the problem still remains in the ascii output. This is bug 7057: summation is ambiguous.

Contents

2

— * —

)set break resume
)spool summation.output
)set message test on
)set output mathml on
)set message auto off
)clear all
--S 1 of 5
summation(i^2,i=a..b)^(d-c)
--R
--R         b      d - c
--R        --+    2
--R   (1)  >     i
--R        --+
--R        i= a
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><msup><mrow><mo>(</mo><mrow><munderover><mo>&#x02211;</mo>...
--R</math>
--R
--R                                                    Type: Expression(Integer)
--E 1
--S 2 of 5
summation(i^2^(d-c),i=a..b)
--R
--R         b      d - c
--R        --+    2
--R   (2)  >     i
--R        --+
--R        i= a
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><munderover><mo>&#x02211;</mo><mrow><mrow><mrow><mi>i</mi>...
--R</math>
--R
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 5
sum(operator(f) (i)+1,i=1..n)
--R
--R         n
--R        --+
--R   (3)  >     f(i) + 1
--R        --+
--R        i= 1
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><munderover><mo>&#x02211;</mo><mrow><mrow><mrow><mi>i</mi>...
--R</math>
--R

--R                                                    Type: Expression(Integer)
--E 3
--S 4 of 5
sum(operator(f) (i),i=1..n)+1
--R
--R         n
--R        --+
--R   (4)  >     f(i) + 1
--R        --+
--R        i= 1
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><mo>(</mo><mrow><munderover><mo>&#x02211;</mo><mrow><mrow>...
--R</math>
--R
--R                                                    Type: Expression(Integer)
--E 4
--S 5 of 5
sum(operator(f) (i)+1,i=1..n)^3
--R
--R         n            3
--R        --+
--R   (5)  >     f(i) + 1
--R        --+
--R        i= 1
--R<math xmlns="http://www.w3.org/1998/Math/MathML" mathsize="big" display="block">
--I<mrow><msup><mrow><mo>(</mo><mrow><munderover><mo>&#x02211;</mo>...
--R</math>
--R
--R                                                    Type: Expression(Integer)
--E 5
)spool
)lisp (bye)

———-

References^{[1] nothing}

5